OpenStudy (mew55):

integral 0 to 1 ln(x)dx

OpenStudy (turingtest):

integrate by parts du=dx, u=lnx

OpenStudy (turingtest):

dv=dx, u=lnx I mean...

OpenStudy (mew55):

so just taking the derivative of ln(x) but isnt it 1/x?

OpenStudy (turingtest):

yep

OpenStudy (mew55):

then after getting the derivtive just put the 0 to 1 limit?

OpenStudy (turingtest):

what? I said integrate by parts, do you know how to do that?

OpenStudy (zepp):

Um, \[\int \ln(x)dx\ne\frac{1}{x}+C\] D:

OpenStudy (turingtest):

@ivanmlerner I hope you're not typing the full solution

OpenStudy (anonymous):

No, relax

OpenStudy (mew55):

but for imporoper integrals, we have to to take the limits

OpenStudy (turingtest):

yes, but first you gotta integrate it, the limits come in at the evaluation

OpenStudy (anonymous):

im trying to explain the integration by parts from scratch

OpenStudy (turingtest):

okay, don't let me stop you. sorry to interrupt but I didn't want you to jump the gun.

OpenStudy (mew55):

im not asking or answers, i am just trying to understand how to set it up :(

OpenStudy (anonymous):

Integration by parts is the analogous (not sure if this word exists in english) of the product rule for derivatives. Product rule: \[\frac{ d(f(x)g(x)) }{ dx }=\frac{ df }{ dx }g+\frac{ dg }{ dx }f\] Integrating that in x, we get: \[f(x)g(x)=\int\limits_{}^{}\frac{ df }{ dx }g dx+\int\limits_{}^{}\frac{ dg }{ dx }f dx\] When you put one of the integrals to the other side you get: \[f(x)g(x)-\int\limits\limits_{}^{}\frac{ df }{ dx }g dx=\int\limits\limits_{}^{}\frac{ dg }{ dx }f dx\] Now the only problem, is to find g and f so that it becomes simpler to solve.

OpenStudy (mew55):

its \[\int\limits_{1}^{0}lnxdx\]

OpenStudy (turingtest):

so use what @ivanmlerner showed you and sub in u=lnx dv=dx

OpenStudy (turingtest):

the way I write integration by pars, which is the same as ivans, is\[\int udv=uv-\int vdu\]

OpenStudy (mew55):

i was asking someone else XD but thankx. i'll try the integration by parts

OpenStudy (turingtest):

sorry, I didn't see you ask anything :/